# TimeScope: Towards Task-Oriented Temporal Grounding In Long Videos

### v2

@@ 4.1 Implementation details
| TimeChat Ren et al. (2024) | 32.2 | 13.4 | 32.2 | 36.2 | 20.2 | 21.8 |
| VTimeLLM Huang et al. (2024a) | 27.5 | 11.4 | 31.2 | 44.0 | 27.8 | 30.4 |
| VideoChat-Flash Li et al. (2024) | 53.1 | 27.6 | – | – | – | – |
| TRACE Guo et al. (2024) | 61.7 | 41.4 | 41.4 | 37.7 | 24.0 | 39.0 |
| HawkEye Wang et al. (2024) | 58.3 | 28.8 | – | 55.9 | 34.7 | – |
| TimeSuite Zeng et al. (2025) | 67.1 | 43.0 | – | – | – | – |
| Time-R1 Wang et al. (2025b) | 72.2 | 50.1 | – | 58.6 | 39.0 | – |
| DeepVideo-R1-7B Park et al. (2025) | 71.7 | 50.6 | 61.2 | 33.9 | 18.0 | 36.9 |
| VideoChat-R1-7B Li et al. (2025c) | 71.7 | 50.2 | 60.8 | 33.4 | 17.7 | 36.6 |
| TimeZero-7B Wang et al. (2025b) | 60.8 | 35.3 | 58.1 | 39.0 | 21.4 | 40.5 |
| Temporal-RLT-7BLi et al. (2025a) | 67.9 | 44.1 | 57.0 | 38.4 | 20.2 | 39.0 |
| TimeScope-7B | 78.9 | 64.0 | 57.5 | 69.6 | 59.0 | 48.0 |
4.2 Benchmarks and Metrics
For evaluation, we benchmark our method across the following four categories: (i) short-video temporal grounding, including Charades-STA [46] and ActivityNet-Captions; (ii) long-video temporal grounding, including videos longer than 300s from V-STaR and Vid-Chapters-7M; (iii) Video Question Answering (VideoQA), including three general long-video QA benchmarks: CG-Bench, MLVU, and LongVideoBench ; (iv) long video task grounding, including our proposed ToTG-bench. The statistics of the evaluation benchmarks used are listed in Table 5.
Evaluation metrics. For temporal grounding tasks and task grounding tasks, we adopt Recall@1 (R@1) at multiple temporal intersection-over-union (IoU) thresholds and mean IoU (mIoU) as evaluation metrics. Specifically, for temporal grounding benchmarks, we use IoU thresholds of 0.3, 0.5, and 0.7. For general VideoQA tasks, we report standard accuracy metrics.
| Model | R1@0.5 | R1@0.7 |
| VTimeLLM-7B Huang et al. (2024a) | 4.1 | 1.6 |
| CLIP Radford et al. (2021) | 5.2 | 2.3 |
| M-DETR Kamath et al. (2021) | 27.3 | 17.6 |
| ReVisionLLM Hannan et al. (2024) | 27.4 | 21.8 |
| Qwen2.5-VL-7B Bai et al. (2025) | 0.3 | 0.1 |
| VITAL-7B Zhang et al. (2025) | 25.8 | 19.5 |
| TimeScope-7B | 24.3 | 22.1 |
@@ 4.2 Benchmarks and Metrics
| UniTime | 0.629 | 0.629 |
| Keye-1.5-VL-8B | 0.639 | 0.491 |
| TimeScope-7B | 0.909 | 0.909 |
@@
| UniVTG Lin et al. (2023) | 34.87 | 62.56 | 54.67 |
| VTimeLLM Huang et al. (2024a) | 34.60 | 59.52 | 54.30 |
| TimeSuite Zeng et al. (2025) | 32.47 | 58.51 | 53.25 |
| UniTime-Full Li et al. (2025f) | 40.30 | 66.50 | 56.47 |
| TimeScope-7B | 39.53 | 70.14 | 59.18 |
| Method | S(R1@0.5) | S(R1@0.7) | M(R1@0.5) | M(R1@0.7) | L(R1@0.5) | L(R1@0.7) |
| no option | ||||||
| Qwen-2.5VL-7B | 0.458 | 0.400 | 0.111 | 0.091 | 0.030 | 0.030 |
| ARC-Hunyuan-Video-7B | 0.523 | 0.484 | 0.36 | 0.260 | 0.175 | 0.162 |
| Keye-VL-1.5-8B | 0.521 | 0.333 | 0.313 | 0.261 | 0.221 | 0.124 |
| UniTime | 0.524 | 0.427 | 0.296 | 0.232 | 0.242 | 0.226 |
| TimeScope-7B | 0.574 | 0.436 | 0.47 | 0.44 | 0.435 | 0.405 |
| has option | ||||||
| Qwen-2.5VL-7B | 0.475 | 0.413 | 0.111 | 0.091 | 0.030 | 0.030 |
| ARC-Hunyuan-Video-7B | 0.158 | 0.148 | 0.111 | 0.090 | 0.124 | 0.092 |
| Keye-1.5-VL-8B | 0.527 | 0.363 | 0.514 | 0.471 | 0.261 | 0.185 |
| UniTime | 0.557 | 0.441 | 0.317 | 0.274 | 0.287 | 0.256 |
| TimeScope-7B | 0.683 | 0.515 | 0.52 | 0.44 | 0.467 | 0.426 |
4.3 SOTA Performance on Traditional Temporal Grounding
We conduct a comprehensive comparison of TimeScope against dual-encoder-based methods, MLLM-based methods, and video understanding models on both short- and long-video temporal grounding benchmarks. As shown in Table 1, TimeScope achieves state-of-the-art (SOTA) performance across all benchmarks. For instance, on Charades-STA, TimeScope reached an R1@0.7 score of 64.0, surpassing VideoChat-Flash (27.6), TimeSuite (43.0), and Time-R1 (50.1). On ActivityNet, it achieved an R1@0.7 score of 59, outperforming HawkEye (34.7) and Time-R1 (39.0).
Particularly noteworthy is that, compared to most other models, TimeScope maintains a smaller gap between R@1@0.5 and R@1@0.7. This indicates that our model is capable of more precise temporal localization. Furthermore, as shown in Table 3, TimeScope achieved an R1@0.7 score of 90.9 on V-STaR, where video duration reaches 300 seconds, significantly surpassing UniTime’s 62.9 and Keye-1.5-VL’s 49.1, demonstrating its superior performance and strong capability in handling long-video grounding tasks.
4.4 SOTA Performance on Task-oriented temporal grounding
Besides traditional temporal grounding, we also evaluate TimeScope on ToTG-Bench, comparing it with MLLM-based models and generic MLLMs. As shown in Table LABEL:lvtgbench, where “has option” indicates that the options are fed into the model along with certain prompt information, TimeScope achieves outstanding and robust performance across different video durations. Specifically, while TimeScope and ARC-Hunyuan-Video-7B Ge et al. (2025) achieve comparable performance on short videos (within a 5-point difference), TimeScope shows clear superiority on medium and long videos, where it outperforms other models by 10 to 20 points. This indicates that our approach can effectively handle videos of various durations, especially long videos, and demonstrates excellent grounding ability for task-oriented queries.
| Model | Early | Middle | Late | Std Dev () |
| Qwen-2.5-VL Bai et al. (2025) | 20.0 | 13.4 | 4.1 | 8.0 |
| Keye-1.5-VL Yang et al. (2025a) | 50.0 | 33.3 | 31.1 | 10.4 |
| UniTime-Full Yang et al. (2025a) | 50.9 | 28.1 | 44.0 | 11.6 |
| TimeScope-7B | 39.3 | 44.8 | 46.0 | 3.5 |
4.5 Improvement for Long-Video QA
As discussed in Section 3.1, the strong performance of TimeScope on the ToTG task suggests that it can help MLLMs capture the critical information in long videos required for question answering. To validate this point, we first use TimeScope and other temporal grounding models to localize time interval. The frames within the predicted intervals are then fed into Qwen2.5-VL-7B for answering. Their results are compared against the default setting, where frames are uniformly sampled without temporal grounding. We evaluate this experiment on three long-video understanding benchmarks: CG-Bench, MLVU, and LongVideoBench. As shown in Table 4(a), TimeScope consistently brings significant improvements across all benchmarks, while other temporal grounding models fail to outperform the default uniform-sampling baseline.
| Model | R1@0.3 | R1@0.5 | R1@0.7 |
| TimeScopeScope | |||
| Zeroshot | 89.7 | 85.2 | 73.8 |
| TimeScopeScope | |||
| Multi-stage | 92.1 | 90.9 | 90.9 |
| Method | Prefill | Decode | Cost |
| zero shot | |||
| Stage1 | 1400ms | 672ms | 2072ms |
| Multi-stage | |||
| Stage1 | 1400ms | 441ms | |
| Stage2 | 158ms | 357ms | 2356ms |
@@ 4.6 Ablation Studies
Effectiveness of Progressive Reasoning. To assess the effectiveness and necessity of progressive reasoning, we conduct an ablation study comparing two settings: progressive reasoning and single-step reasoning. The study is performed on videos longer than 300 seconds from the V-STaR benchmark, with results reported in Table 4(a). We observe that progressive reasoning achieves substantial improvements over the single-step baseline. In particular, for longer videos, progressively narrowing the search space proves both effective and necessary for accurate temporal grounding.
Efficiency of Progressive Reasoning. In addition to effectiveness, we also evaluate the computational efficiency of TimeScope. As shown in Table 4(b), progressive reasoning introduces only marginal additional overhead compared to single-step reasoning, while still delivering substantial accuracy gains. This efficiency largely benefits from our design of adjusting KV sparsity during progressive reasoning. Overall, TimeScope strikes a cost-effective balance between accuracy and computational cost, making it practical for long-video temporal grounding.
Robustness against Time Bias. After analyzing the evaluation results on ToTG-Bench, we find that the Query Center (the position of the target time interval) introduces a significant bias that strongly influences the performance of temporal grounding models. As shown in Table 6, Qwen2.5-VL and Keye-1.5-VL tend to perform better when the target time interval appears at the beginning of the video, but their performance drops substantially when the target lies in the middle or towards the end. In contrast, our TimeScope model maintains robust performance across different target interval positions, demonstrating its resilience against this bias.
@@ 5 Conclusion
Appendix A Use of LLMs
In the process of writing this paper, we utilized a Large Language Model (LLM) solely as a tool for polishing the language and enhancing the readability of the manuscript. The LLM was employed to refine the wording, grammar, and coherence of the text, ensuring that the content is clear and well-structured. However, it is important to note that the LLM did not play any significant role in the conception of research ideas or the formulation of the paper’s content. All research ideas, methodologies, and conclusions presented in this paper are the original work of the authors. We understand that we bear full responsibility for the contents of this paper, including any text that has been processed by the LLM. We have carefully reviewed and verified all content to ensure its accuracy and originality. Any contributions made by the LLM were limited to language enhancement and did not involve any form of plagiarism or scientific misconduct.
Appendix B Benchmark Details

### v3

@@ 4.2 TimeScope
| TimeChat [32] | 32.2 | 13.4 | 32.2 | 36.2 | 20.2 | 21.8 |
| VTimeLLM [13] | 27.5 | 11.4 | 31.2 | 44.0 | 27.8 | 30.4 |
| VideoChat-Flash [19] | 53.1 | 27.6 | – | – | – | – |
| TRACE [12] | 61.7 | 41.4 | 41.4 | 37.7 | 24.0 | 39.0 |
| HawkEye [37] | 58.3 | 28.8 | – | 55.9 | 34.7 | – |
| TimeSuite [44] | 67.1 | 43.0 | – | – | – | – |
| Time-R1 [38] | 72.2 | 50.1 | – | 58.6 | 39.0 | – |
| DeepVideo-R1-7B [28] | 71.7 | 50.6 | 61.2 | 33.9 | 18.0 | 36.9 |
| VideoChat-R1-7B [21] | 71.7 | 50.2 | 60.8 | 33.4 | 17.7 | 36.6 |
| TimeZero-7B [38] | 60.8 | 35.3 | 58.1 | 39.0 | 21.4 | 40.5 |
| Temporal-RLT-7B[18] | 67.9 | 44.1 | 57.0 | 38.4 | 20.2 | 39.0 |
| TimeScope-7B | 78.9 | 61.2 | 56.2 | 66.9 | 56.0 | 46.0 |
Progressive Reasoning. With both and prepared, TimeScope performs grounding in a coarse-to-fine manner. In the first step, only the compact holistic representation is kept in GPU memory, while the high-resolution remains stored in CPU memory. Leveraging the lightweight long-range context encoded in , the MLLM efficiently predicts a coarse temporal interval that is most likely to contain the task-relevant moment for the query . This stage effectively eliminates the majority of irrelevant frames and substantially reduces the temporal search space. Next, TimeScope reloads only the fine-grained corresponding to the predicted interval and performs detailed reasoning. The rich local temporal and visual details preserved in allow the model to refine the boundaries within the coarse region and accurately localize the target moment. By combining the global efficiency of with the local precision of , the progressive reasoning process achieves high localization accuracy while maintaining low computational cost of processing the long video at high resolution.
4.3 ToTG-Pile
To maximize TimeScope’s capacity for task-oriented temporal grounding, we created ToTG-Pile—a large-scale dataset bridging conventional temporal grounding data with Task-oriented Temporal Grounding requirements. ToTG-Pile features two key characteristics: (1) all samples follow task-oriented query definitions (Sec. 3), where queries are formulated as task instructions rather than explicit temporal descriptions, and (2) each sample includes both ground-truth temporal intervals and chain-of-thought (CoT) annotations that detail the localization reasoning process.
We constructed the dataset by collecting diverse videos with broad coverage of tasks and visual contexts. Specifically, we leverage comprehensive Video VQA datasets such as VideoR1 [9] and FineVideo [8]. To curate high-quality reasoning data, we employ a three-stage pipeline: (1) Answer-aware prompting: we feed answers from original VQA samples into expert temporal grounding models to obtain candidate temporal intervals; (2) Cross-validation filtering: we leverage multiple expert models for cross-validation, discarding low-quality samples with IoU 0.1 across models; (3) CoT annotation generation: we use reasoning-capable MLLMs to generate chain-of-thought annotations that detail the localization reasoning process. Additionally, to extend the dataset to long-video scenarios, we synthesized 90K long clips (10 minutes each) by concatenating short videos and generated corresponding temporal grounding queries from video captions. This augmentation enhances TimeScope’s capability in handling both conventional temporal grounding and task-oriented temporal grounding in extended video contexts.
Overall, ToTG-Pile provides a large-scale, diverse, and reasoning-oriented foundation for training MLLMs to perform effective and generalizable task-oriented temporal grounding in long videos.
5 Experiment
5.1 Implementation details
We adopt VideoXL-2 [29] as our backbone for two reasons: (1) it can process very long video sequences, enabling straightforward construction of long-video temporal understanding methods, and (2) its internal design interleaves timestamp tokens, providing the model with a strong built-in temporal awareness. Stage 1: Basic Localization. We use the temporal grounding splits of ToTG-Pile and train the model to predict target time intervals directly from raw video and task descriptions, bootstrapping its basic localization ability. Stage 2: Coarse-to-Fine Refinement. We apply heavy temporal augmentations (random cropping, shifting, and scaling of time spans) to training videos, forcing the model to first estimate a coarse temporal window from abstract video representations and then refine it into fine-grained intervals using detailed representations.
5.2 Results on Task-oriented temporal grounding
Evaluation on ToTG-Bench. We evaluate TimeScope on ToTG-Bench, which categorizes videos into three duration segments: short (180s), medium (180–600s), and long (600s). We report IoU@0.3 and IoU@0.5 metrics and compare TimeScope against both specialized temporal understanding MLLMs and general video understanding models. As shown in Table 2, TimeScope demonstrates strong and consistent performance across all video durations. Most notably, TimeScope outperforms all baselines by 20–30 points on medium and long videos, demonstrating substantial advantages in handling task-oriented temporal grounding in extended contexts. On short videos, TimeScope achieves competitive performance, trailing the state-of-the-art Temporal-RLT [18] by less than 2 points while significantly surpassing it on longer videos. These results highlight TimeScope’s effectiveness in task-oriented temporal localization, particularly its ability to perform accurate reasoning and grounding in long-video scenarios while maintaining robust performance across different video durations.
We further evaluate TimeScope and baseline models on ToTG-Bench across different task categories using IoU@0.5 metric, as shown in Table 3. TimeScope demonstrates strong performance across all task categories. Notably, in reasoning tasks (including Object Reasoning, Spatial Reasoning, Temporal Reasoning, and Action Reasoning), TimeScope outperforms all baselines by 10–20 points, highlighting its superior capability in task-oriented reasoning and temporal localization. Across other task categories, TimeScope consistently achieves top-tier performance, demonstrating robust generalization to diverse task types.
@@ 5.2 Results on Task-oriented temporal grounding
| UniTime | 62.9 | 62.9 |
| Keye-1.5-VL-8B | 63.9 | 49.1 |
| TimeScope-7B | 87.5 | 85.2 |
5.3 Results on Traditional Temporal Grounding
We conduct a comprehensive comparison of TimeScope against traditional and MLLM-based methods on conventional temporal grounding benchmarks, covering both short-video and long-video settings.
Short-video Benchmarks. As shown in Table 4, TimeScope achieves state-of-the-art performance across all short-video benchmarks. On Charades-STA, TimeScope attains an R1@0.7 score of 64.0, significantly surpassing VideoChat-Flash (27.6), TimeSuite (43.0), and Time-R1 (50.1). On ActivityNet, it achieves an R1@0.7 score of 59.0, outperforming HawkEye (34.7) and Time-R1 (39.0). Notably, TimeScope maintains a smaller gap between R1@0.5 and R1@0.7 compared to most baselines, indicating its capability for more precise temporal localization.
Long-video Benchmarks. As shown in Table 5, TimeScope achieves an R1@0.7 score of 85.2 on V-STaR (with videos up to 300 seconds), substantially exceeding UniTime (62.9) and Keye-1.5-VL (49.1). This demonstrates TimeScope’s strong capability in handling extended temporal grounding tasks.
@@ 5.3 Results on Traditional Temporal Grounding
| UniVTG [24] | 34.87 | 62.56 | 54.67 |
| VTimeLLM [13] | 34.60 | 59.52 | 54.30 |
| TimeSuite [44] | 32.47 | 58.51 | 53.25 |
| UniTime-Full [23] | 40.30 | 66.50 | 56.47 |
| TimeScope-7B | 38.47 | 68.12 | 58.34 |
5.4 Benefits to Long-Video Understanding
As discussed in Section 3, TimeScope’s strong performance on task-oriented temporal grounding demonstrates its potential to help MLLMs capture critical information in long videos for question answering. To validate this, we conduct experiments where TimeScope and baseline temporal grounding models first localize relevant time intervals, and then feed frames from the predicted intervals into Qwen2-VL-7B [1] for answer generation. We compare these results against other grounding models and a default uniform sampling baseline without temporal grounding.
We evaluate on three long-video understanding benchmarks: CG-Bench, MLVU, and LongVideoBench. As shown in Table 6, TimeScope demonstrates strong performance and brings substantial improvements over uniform sampling across all benchmarks, surpassing most temporal grounding baselines. On MLVU and LongVideoBench, TimeScope achieves the highest scores of 68.12 and 58.34 respectively, demonstrating its effectiveness in identifying task-relevant temporal segments for long-video understanding. Notably, while many video understanding questions require information beyond a single temporal segment, TimeScope’s task-oriented grounding still provides meaningful performance gains, validating its practical utility for downstream applications.
@@ 5.5 Ablation Studies
Effectiveness of Progressive Reasoning. To evaluate the effectiveness and necessity of progressive reasoning in long-video scenarios, we conducted an ablation study comparing two settings: progressive reasoning versus standard single prediction. The experiment was performed on samples longer than 300 seconds from the V-STaR benchmark and samples exceeding 600 seconds from the ToTG-Bench, with results presented in Table 7. The findings demonstrate that progressive reasoning achieves substantial improvements over the single-step baseline across both traditional temporal grounding tasks and task-oriented tasks in long videos. Particularly notable is the performance gain huge improvement in IoU@0.7, which confirms the effectiveness of progressively narrowing the search space for achieving precise temporal localization.
Efficiency of Progressive Reasoning. Despite employing progressive reasoning to achieve higher prediction accuracy, TimeScope remains highly efficient—thanks to its streaming-based video representation construction and carefully designed dual-level representation. As shown in Table 8, our framework enables the MLLM backbone to process significantly more frames while simultaneously improving throughput in both the prefill and decode stages, outperforming baseline methods without TimeScope in both speed and scalability.
| Method | V-STaR(long) | ToTG-bench(long) | ||
| R1@0.5 | R1@0.7 | R1@0.5 | R1@0.7 | |
| Zeroshot | 73.2 | 61.4 | 37.8 | 29.5 |
| Progressive Reasoning | 86.4 | 83.0 | 37.8 | 34.0 |
Effectiveness of Holistic Video Representation. As outlined in Sec. 4.2, TimeScope enables dense frame sampling for long videos—made feasible by the efficient design of its Holistic Video Representation (HVR). As demonstrated in Table 9(a), with HVR, TimeScope supports dense sampling even for extremely long videos (up to 2,000 frames). This capability directly translates into significant gains in temporal grounding precision, particularly for fine-grained event localization.
| Method | Frames | Prefill | Decode | Sum |
| w/o TimeScope | 800 | 2903ms | 816ms | 3719ms |
| w TimeScope | 800 | 2259ms | 517ms | 2776ms |
| 1200 | 3205ms | 627ms | 3832ms |
Efficiency of ToTG-pile Dataset. To further analyze the impact of the ToTG-Pile dataset, we retrained the model by excluding it from the training set and reported the corresponding results on short-video temporal benchmark Charades-STA, long-video benchmarks V-STaR and ToTG-bench (long), as shown in Table 9(b). It can be observed that the performance gap becomes particularly pronounced on V-STaR and ToTG-bench (long), where ToTG-pile brings performance gains of 40 and 20 points, respectively, compared to the setting without ToTG-pile. This demonstrates the crucial value of ToTG-pile in establishing TimeScope’s long-video understanding capability and task-oriented reasoning proficiency.
| Setting | Max Frames | V-STaR(long) | ToTG-Bench(long) | ||
| R1@0.5 | R1@0.7 | R1@0.5 | R1@0.7 | ||
| w/o Holistic VR. | 800 | 73.2 | 61.4 | 37.8 | 29.5 |
| w Holistic VR. | 2000 | 80.7 | 76.1 | 41.0 | 37.1 |
| Method | Charades-STA | V-STaR(long) | ToTG-bench(long) | |||
| R1@0.5 | R1@0.7 | R1@0.5 | R1@0.7 | R1@0.5 | R1@0.7 | |
| W/o ToTG-Pile | 74.9 | 55.8 | 47.7 | 39.7 | 29.4 | 22.7 |
| W ToTG-Pile | 78.9 | 61.2 | 86.4 | 83.0 | 37.8 | 34.0 |
6 Qualitative results
We show the results of TimeScope in Figure 3, which can be seen that TimeScope exhibits robust performance with good coarse-grained segment retrieval and fine-grained temporal grounding capabilities. More results can be seen in Appendix.
@@ 7 Conclusion
Appendix A Different and Other Grounding-QA
Although some traditional groundingQA datasets—such as CG-Bench, NExT-GQA, and ReXTime already require localizing temporal evidence based on the question rather than on explicit event phrases, the Task-oriented Temporal Grounding task we propose is fundamentally different. These differences are illustrated in Figure 1. Specifically, questions in conventional question-based grounding works usually still contain words strongly correlated with the target time interval, allowing existing temporal-grounding models to localize the interval by relying on the dominant cues in the question. In contrast, our task-oriented setting emphasizes reasoning and thinking driven by the question: our questions typically lack descriptors that are directly tied to the answer, and instead demand that the model first infer which temporal segments can solve the question and then ground them.
Appendix B Holistic Video Representation Detail
We conducted a comprehensive ablation on the parameters of Holistic Video Representation (HVR), including the chunk size used to partition the video and the compression ratio applied when generating HVR. We obtain HVR from Detail Video Representation through a pooling operation. During training, we randomly select one-third of the samples to be trained with HVR. As shown in Tables 1 and 2, although using HVR leads to a slight performance drop when the video is input at 1 fps with a maximum of 800 frames, we observe clear performance gains when the maximum number of input frames is increased to 2,000. With HVR, the model surpasses the original zeroshot results.
| Method | Max Frames | V-STaR(long) | ToTG-Bench(long) | |||
| VideoChunk | HVR Rate | R1@0.5 | R1@0.7 | R1@0.5 | R1@0.7 | |
| 0 | 1* | 800 | 73.2 | 61.4 | 37.8 | 29.5 |
| 0 | 2* | 800 | 63.6 | 63.6 | 31.2 | 27.7 |
| 0 | 4* | 800 | 62.5 | 60.2 | 30.4 | 26.4 |
| 150 | 1* | 800 | 61.3 | 60.2 | 29.5 | 26.1 |
| 150 | 2* | 800 | 63.6 | 63.6 | 31.8 | 27.3 |
| 150 | 4* | 800 | 62.5 | 61.3 | 30.6 | 26.9 |
| 300 | 1* | 800 | 60.2 | 58.0 | 28.7 | 24.9 |
| 300 | 2* | 800 | 63.6 | 63.6 | 32.4 | 26.3 |
| 300 | 4* | 800 | 61.3 | 59.1 | 28.6 | 25.8 |
| Method | Max Frames | V-STaR(long) | ToTG-Bench(long) | |||
| VideoChunk | HVR Rate | R1@0.5 | R1@0.7 | R1@0.5 | R1@0.7 | |
| 0 | 1* | 800 | 73.2 | 61.4 | 37.8 | 29.5 |
| 0 | 1* | 2000 | - | - | - | - |
| 150 | 1* | 2000 | 62.5 | 60.2 | 31.4 | 26.7 |
| 150 | 2* | 2000 | 80.7 | 76.1 | 41.0 | 37.1 |
| 150 | 4* | 2000 | 86.4 | 81.8 | 43.2 | 39.7 |
| 300 | 1* | 2000 | 64.7 | 62.5 | 33.4 | 28.2 |
| 300 | 2* | 2000 | 79.5 | 78.4 | 40.6 | 37.3 |
| 300 | 4* | 2000 | 81.8 | 80.6 | 42.7 | 38.8 |
Appendix C ToTG-bench detail
ToTG-Bench comprises 337 videos and nearly 500 questions. As shown in Figure 2, it exhibits rich diversity in task types, video categories, video durations, and the temporal locations of target intervals. The average video length is 805 seconds. Specifically, the benchmark covers 13 task types—such as action reasoning, OCR perception, and temporal reasoning—spanning a wide range of video-understanding tasks.111We adopt the definitions of task type and video category from Video-MME. In addition, it includes 35 video categories from numerous real-world domains. Video durations range from a few minutes to one hour, and the target intervals are uniformly distributed along the timeline to ensure an unbiased evaluation.
Explanations of the 13 Task Categories in ToTG-Bench:
OR (Object Reasoning): Identify why an object is present and how it causally affects later events in the clip.
SR (Spatial Reasoning): Infer spatial relationships (e.g., behind, left-of) between objects that are never simultaneously visible.
TR (Temporal Reasoning): Determine the correct chronological order of events shown out of sequence or with missing segments.
AR (Action Reasoning): Predict the next action an agent will perform given the current context and goals.
SP (Spatial Perception): Locate and describe where objects are in the 3-D scene from egocentric or exocentric views.
AP (Attribute Perception): Recognize object properties such as color, material, or state that remain constant across frames.
TP (Temporal Perception): Detect when a change or event occurs and report its exact start and end times.
EG (Ego): Understand what the camera-wearer is doing, attending to, or will do next from first-person video.
OP (OCR Problems): Read and transcribe text that appears on screens, signs, or documents within the video.
AT (Action Recognition): Classify which predefined action is being performed in a given segment of the video.
CP (Counting Problem): Accurately count how many instances of a specified object or person appear throughout the clip.
TU (Tutorial): Follow instructional video content and answer questions about the steps or techniques demonstrated.
Appendix D Metric